OpenStudy (anonymous):

simplify 121^1/6 121^5/6 ---------------- 121^18/12

OpenStudy (anonymous):



OpenStudy (anonymous):

what is \(\frac{1}{6}+\frac{5}{6}\) ?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

k so the numerators is \(121^1=121\)

OpenStudy (anonymous):

what is \(\frac{18}{12}\) reduced?

OpenStudy (anonymous):

9/6

OpenStudy (anonymous):

reduced

OpenStudy (anonymous):

3/2